# Breaking Down and Building Up: Mixture of Skill-Based Vision-and-Language Navigation Agents

### v1

@@ Introduction
Table 3: Evaluation of each skill-based agent on the NavNuances benchmark across four skill categories: Direction Change (DC), Vertical Movement (VM), Landmark Recognition (LR), and Room Recognition (RR). Ident.: Identification.
Evaluation Datasets.
We primarily use the Room-to-Room (R2R) dataset (Anderson et al. 2018b), especially on its unseen split of validation (Val Unseen) and test (Test Unseen) datasets. R2R is a standard benchmark in VLN consisting of panoramic RGB-D scans from the Matterport3D (Chang et al. 2017) simulator and providing crowd-sourced instructions paired with paths. Besides, we evaluate the generalization ability of SkillNav on GSA-R2R (Hong et al. 2025) which includes residential (R) and non-residential (N) scenes (e.g. shops, restaurants, and museums) from Habitat-Matterport3D (Ramakrishnan et al. 2021), and diverse instruction styles with role-specific dialogues (e.g. travel guides (Scene) beyond the basic style of R2R (Basic).
Evaluation Metrics.
We use the standard metrics to evaluate the navigation performance (Anderson et al. 2018a; Zhao, Qi, and Wu 2023): (1) Navigation Error (NE): the distance between the stop location and the target; (2) Oracle Success Rate (OSR): the agent ever gets close enough to the goal at any point along its trajectory, regardless of where it decides to stop; (3) Success Rate (SR): the ratio of agents stopping within $3$ meters of the target; (4) Success rate weighted by Path Length (SPL): measure navigation efficiency by weighting the success rate with the ratio between the shortest path length and the agent’s actual path length, penalizing unnecessarily long trajectories.
Implementation Details.
We utilize CLIP-B/16 (Radford et al. 2021) as the visual backbone and BERT-base-uncased (Devlin et al. 2018) as the language backbone within our DUET-based skill agents. During the skill training, we fine-tune the DUET pre-trained model with Temporal Order synthetic data, ScaleVLN augmentation data, and R2R Train data for $50,000$ iterations using a batch size of 32 and a learning rate of $5\times 10^{-5}$ on 1 NVIDIA A6000 GPU with the random seed 0. The best finetuned Temporal DUET model is selected based on the SPL performance on the R2R Validation Unseen dataset. Based on the Temporal DUET, we employ the second round fine-tuning with atomic skill synthetic data for 30,000 iterations with a batch size of 16 on the same GPU. In our SkillNav LLM-based architecture, we adopt GPT-4o (OpenAI 2024) as the Temporal Reordering module due to its superior instruction-following capabilities and employ Qwen2.5-VL-7B-Instruct (Bai et al. 2025) as the action router because of its strong multi-modal alignment and reasoning abilities. All inferences with the action router are performed using in-context prompting. More details are provided in the supplementary material.
@@
As shown in Table 2, SkillNav achieves strong overall performance across both R2R datasets and demonstrates robust generalization on GSA-R2R, outperforming most fine-tuned and LLM-based agents. On the R2R unseen environments, SkillNav (Method #11) achieves $83\%$ SR and $77\%$ SPL, achieving second overall, behind only SRDF (Method #10). While SRDF achieves the highest performance on R2R Test-Unseen, this can be largely attributed to its pretraining on large-scale data that closely follows R2R-style instruction patterns. However, this reliance limits its generalization ability on GSA-R2R, with $13\%$ and $5\%$ SR drop on Test-N-Basic and Test-N-Scene, respectively. This highlights the importance and efficacy of explicitly learning atomic navigation skills. SkillNav outperforms another strong baseline, ScaleVLN (Method #9), by $4\%$ in SR and notably by $7\%$ in SPL, indicating that SkillNav predicts more efficient and precise navigation trajectories. Additionally, SkillNav also demonstrates SOTA generalization performance in GSA-R2R, ranking 1st in SPL across all GSA-R2R splits. Notably, on Test-N-Scene, which combines non-residential environments with more complex and role-specific instructions, SkillNav matches the best SR tied with NavGPT-2 (Method #3), while significantly outperforming it in SPL. NavGPT-2 benefits from fine-tuning on FlanT5-XXL (Chung et al. 2022), which likely enhances its ability to interpret stylized instructions. However, its lower SPL reveals inefficiencies in path planning and execution. While LLMs can help parse diverse instruction styles, they often introduce noise or lose critical spatial details when translating, limiting their effectiveness in downstream navigation tasks. This highlights the need for tightly integrated skill reasoning and grounded visual understanding, beyond language interpretation alone.
Ablation Study and Analysis
@@
Ablation study on skill-based agents.
To further probe the capabilities of our skill-based agents, we have a fine-grained evaluation on the NavNuances (Wang et al. 2024c), which categorizes navigation instructions into four atomic skills: (1) Direction Change (DC), (2) Vertical Movement (VM), (3) Landmark Recognition (LR), and (4) Region Recognition (RR). These subsets isolate specific reasoning capabilities and allow us to assess each agent’s specialization. As shown in Table 3, each skill-based agent in SkillNav excels in its corresponding category. The Vertical Movement agent achieves the highest SR ($87.65\%$) and SPL ($83.83\%$) on VM, while the Direction Adjustment agent leads in DC with an SR of $70.81\%$. The Landmark Detection agent performs best in LR with $31.53\%$ SR, and the Area and Region Identification agent reaches $85.09\%$ SR on RR. These results validate our skill-based training and data augmentation strategy, confirming that targeted supervision fosters functional specialization that outperforms generalist VLN baselines in isolated skill settings.
Stop Agent Analysis.
@@
Table 4: Ablation results on GSA-R2R across residential (R) and non-residential (N) scenarios with varying instruction styles (Basic and Scene). Reorder: ✔ = LLM-guided Temporal Reordering enabled; ✗ = disabled. Router: Random = randomly select skill-based agents without utilizing action router; Qwen = Qwen2.5-VL-7B-Instruct; GLM = GLM-4.1V-9B-Thinking.
Ablation study on temporal reordering module.
@@
We conduct an ablation study to evaluate of SkillNav’s two key components: the LLM-guided Temporal Reordering module and the VLM-based action router. The results, shown in Table 4, are reported across GSA-R2R splits, covering both residential (R) and non-residential (N) environments with varying instruction styles. First, we evaluate the effectiveness of the temporal reordering module. As shown in rows #2 and #4, when using the same router (Qwen2.5-VL), incorporating the reordering module consistently improves performance across all benchmarks. The improvement is particularly notable in Test-N-Basic, with a significant SPL increase of $1.72\%$, demonstrating that temporally structured subgoals offer clearer guidance for effective skill selection.
Ablation study on action router.
@@
To independently evaluate the effectiveness of our action router, we compare the performance of randomly selected skills without a router (row #1) against our proposed Qwen router. The observed improvements in both SR and SPL metrics clearly indicate the router’s effectiveness: specifically, Test-N-Scene SR increases from $54.61\%$ to $55.46\%$, and SPL rises notably from $43.17\%$ to $45.43\%$. These results confirm that our VLM-based router effectively selects appropriate skills even in the absence of temporal structuring. We further examine the significance of router selection by comparing rows #3 and #4, where the instruction reordering is fixed, and only the router model varies. Qwen2.5-VL-7B-Instruct consistently achieves superior SPL across all splits, particularly notable in Test-N-Scene ($47.96\%$ vs. $46.51\%$), underscoring its enhanced visual grounding capabilities compared to GLM-4.1V-9B-Thinking (Team et al. 2025). This emphasizes that high-quality vision-language representations are essential for effective skill routing in complex and diverse navigation environments.
@@
In Figure 4, we present two qualitative examples highlighting SkillNav’s capability to dynamically select the appropriate skill at each navigation step. These examples illustrate the effectiveness of our approach in reordering temporal action plans, accurately identifying the currently focused subgoal via the router, and subsequently selecting the correct action. Specifically, in Figure 4 (a), the router correctly reasons that the agent has reached the target pillars and decides it is time to stop, resulting in the agent appropriately choosing the stop action at the view containing the pillars. Similarly, in Figure 4 (b), the router identifies the need to move toward the stairs and accordingly selects the vertical movement skill. Overall, SkillNav successfully interprets diverse instruction styles and performs robustly across both residential and non-residential scenes.
@@ Conclusion
To improve the generalization of VLN agents, we introduce SkillNav, a VLN agent that combines skill-based learning with zero-shot VLM-based routing to dynamically select the most suitable actions based on the most relevant expert. We evaluate SkillNav on benchmarks such as R2R and demonstrate its strong generalization capabilities on the challenging GSA-R2R dataset. Our comprehensive experiments further highlight SkillNav’s effectiveness in improving fine-grained skill performance, with detailed ablation studies emphasizing the contributions of the temporal reordering and routing components. We believe our framework provides a novel and interpretable approach that advances compositional reasoning and generalization for the VLN research community.
@@ Acknowledgment
This appendix provides additional details on our method, prompt design, and supporting experiments to complement the main paper, including:
  • Explanations of the synthetic data generation process, including how we construct skill-specific trajectories.
  • Analysis of path length constraints used during trajectory generation to ensure temporal consistency and alignment with existing VLN datasets.
  • Analysis of the effectiveness of the Temporal Order Planning agent and its contribution to resolving failure cases.
  • Full prompt templates used for synthetic instruction generation and agent control.
Primary Factors of Trajectory Generation
As introduced in Section Skill-Specific Data Synthesis and Agent Training in Methodology, we construct 5 skill-specific data and train the agents based on them. The primary factors for the construction of each skill are as follows:
@@
(1) The stop should occur at a place with or after semantically relevant context for pausing, e.g., in front of a painting, at the foot of stairs. (2) The candidate image is very similar with the previous viewpoints.
@@
Path Length in Trajectory Generation
@@
Temporal Order Planning Agent
@@
LLM and VLM Prompts
@@
2
3<Task>:
4Your task is to analyze a list of previously observed images and a natural language instruction.
5Determine which parts of the instruction have already been completed, and return the next step to be executed.
6
7<Response Rules>
8Your response must:
9- Return the next action using *exact phrasing* from the reordered instruction (no paraphrasing).
10- Match the sub-instruction to the visual context from previous images.
11- If the goal (e.g., pool table) has clearly been reached, return the final sub-instruction.
12- If *all* sub-instructions have been completed based on the visual path, do not return anything further. Stop reasoning.
13- If the final destination has been reached and the last step is a positional or waiting action (e.g., ‘‘wait there’’, ‘‘step to the left’’), return that as the next step.
14- You must reason about whether the agent is already at the destination.
15- If the current image shows the goal destination (e.g., inside the room with the pool table, or inside the open doorway), and the instruction contains a final step like ‘‘wait’’ or ‘‘adjust your position’’, that is the next sub-instruction.
16
17---
18
19Use the following reasoning strategy to determine what to do next:
20<Step-by-Step Reasoning Instructions>:
@@
22- Break the full instruction into smaller steps. Each sentence or clause typically represents one step.
23- Example:
@@
30- Do not reissue any previously executed sub-instructions.
31- Compare upcoming steps against what may have been visually completed, even if not explicitly executed one-by-one.
@@
33- Use visual context to infer if *multiple* sub-instructions have been completed in a single transition.
34- If image progression clearly shows the agent has already bypassed an intermediate area or reached a later goal, mark those steps as implicitly complete.
@@
36- If the current image shows the agent at or beyond the target of a sub-instruction, that step can be considered completed.
37- If the current image shows the agent inside the goal location and only a final positional instruction remains (e.g., ‘‘Step slightly to the left’’), return that final instruction.
@@
39- Use exact wording from the original instruction.
40- Do not return instructions that the agent already visually fulfilled, even if they were skipped.
@@
42{
43"Sub-instruction to be executed": "<exact next instruction clause>",
44"Reasoning": "<why this is the next step based on image sequence>"
@@
46CHECKPOINT:
47If multiple sub-instructions were completed based on a single or continuous image segment, skip them and jump to the next logical, visually unfulfilled step.
48
49---
50
51Now, using the instruction and the visual history, identify the next step.
52
53IMPORTANT: Your response must be a valid JSON object without any surrounding text, code blocks, or explanations.
54Do not include markdown formatting like ‘‘‘json or ‘‘‘.
55
56<Original Whole Instruction>:
57‘‘{instruction}’’
58
59<Previous Sub-Instructions>:
60‘‘{previous_sub_instructions}’’
61
62<Previous Viewpoint Images>:
@@
The skill router determines which skill-based agent is most appropriate for executing the selected subgoal among the 5 skill-based agents. Besides, it receives the original instruction I as contextual input to capture additional linguistic cues such as verbs and spatial references. It also uses the reasoning trace from the subgoal localizer to enhance its understanding of the current subgoal. The whole process is displayed in Listing  5.

### v2

@@ Introduction
Evaluation Datasets. We primarily use the Room-to-Room (R2R) dataset (Anderson et al. 2018), especially the unseen split of validation (Val Unseen) and test (Test Unseen) splits. R2R is a commonly-used benchmark in VLN consisting of panoramic RGB-D scans from the Matterport3D (Chang et al. 2017) simulator and providing crowd-sourced instructions paired with navigation paths. Moreover, we evaluate the generalization ability of SkillNav on GSA-R2R (Hong et al. 2025) which includes residential (R) and non-residential (N) scenes (e.g., shops, restaurants, and museums) from Habitat-Matterport3D (Ramakrishnan et al. 2021), and diverse instruction styles with role-specific dialogues (e.g., travel guides (Scene) beyond the basic style of R2R (Basic).
Evaluation Metrics. We use the standard metrics to evaluate the navigation performance (Anderson et al. 2018; Zhao, Qi, and Wu 2023): (1) Navigation Error (NE): the distance between the stop location and the target; (2) Oracle Success Rate (OSR): the agent ever gets close enough to the goal at any point along its trajectory, regardless of where it decides to stop; (3) Success Rate (SR): the ratio of agents stopping within $3$ meters of the target; (4) Success rate weighted by Path Length (SPL): measure navigation efficiency by weighting the success rate with the ratio between the shortest path length and the agent’s actual path length, penalizing unnecessarily long trajectories.
Implementation Details. We utilize CLIP-B/16 (Radford et al. 2021) as the visual backbone and BERT-base-uncased (Devlin et al. 2018) as the language backbone within our DUET-based skill agents. During the skill training, we fine-tune the DUET pre-trained model with Temporal Order synthetic data, ScaleVLN augmentation data, and R2R Train data for $50,000$ iterations using a batch size of $32$ and a learning rate of $5\times 10^{-5}$ on 1 NVIDIA A6000 GPU with the random seed 0. The best finetuned Temporal DUET model is selected based on the SPL performance on the R2R Validation Unseen dataset. Based on the Temporal DUET, we employ the second round fine-tuning with atomic skill synthetic data for $30,000$ iterations with a batch size of $16$ on the same GPU. In our SkillNav LLM-based architecture, we adopt GPT-4o (OpenAI 2024) as the Temporal Reordering module due to its superior instruction-following capabilities and employ Qwen2.5-VL-7B-Instruct (Bai et al. 2025) as the action router because of its strong multi-modal alignment and reasoning abilities. All inferences with the action router are performed using in-context prompting.
@@
As shown in Table 2, SkillNav achieves strong overall performance across both R2R datasets and demonstrates robust generalization on GSA-R2R, outperforming most fine-tuned and LLM-based agents. On the R2R unseen environments, SkillNav (Method #12) achieves $83\%$ SR and $77\%$ SPL, ranking second highest after SRDF (Method #10). While SRDF achieves the highest performance on R2R Test-Unseen, this can be largely attributed to its pretraining on large-scale data that closely follows R2R-style instruction patterns. However, this reliance weakens its generalization ability, leading to a $13\%$ and $5\%$ SR drop on GSA-R2R Test-N-Basic and Test-N-Scene, respectively. SRDF requires additional tuning to remain competitive when transferred to new environments or novel instruction styles. In contrast, SkillNav is trained only on R2R and synthetic skill-specific data, yet achieves strong cross-dataset generalization without any retraining. Additionally, SkillNav also demonstrates SOTA generalization performance in GSA-R2R, ranking 1st in SPL across all GSA-R2R splits and demonstrating its ability to predict more efficient and precise navigation trajectories. Notably, on Test-N-Scene, which combines non-residential environments with more complex and role-specific instructions, SkillNav matches the best SR tied with NavGPT-2 (Method #3), while significantly outperforming it in SPL. NavGPT-2 benefits from fine-tuning on FlanT5-XXL (Chung et al. 2022), which likely enhances its ability to interpret stylized instructions. However, its lower SPL reveals inefficiencies in path planning and execution. While LLMs can help parse diverse instructions, they often introduce noise or lose critical spatial details when translating, limiting their effectiveness in downstream navigation tasks. This highlights the need for tightly integrated skill reasoning and grounded visual understanding, beyond language interpretation alone.
Ablation Study
@@
Skill Evaluation. To further probe the capabilities of our skill-based agents, we have a fine-grained evaluation on the NavNuances, which categorizes navigation instructions into four atomic skills: (1) Direction Change (DC), (2) Vertical Movement (VM), (3) Landmark Recognition (LR), and (4) Region Recognition (RR). These subsets isolate specific reasoning capabilities and allow us to assess each agent’s specialization. As shown in Table 3, each skill-based agent in SkillNav excels in its corresponding category. The Vertical Movement agent achieves the highest SR ($87.65\%$) and SPL ($83.83\%$) on VM, while the Direction Adjustment agent leads in DC with an SR of $70.81\%$. The Landmark Detection agent performs best in LR with $31.53\%$ SR, and the Area and Region Identification agent reaches $85.09\%$ SR on RR. We report the effectiveness of the Stop and Pause agent in Appendix D. These results validate our skill-based training and data augmentation strategy, confirming that targeted supervision fosters functional specialization that outperforms generalist VLN baselines in isolated skill settings.
Table 4: Ablation results on GSA-R2R across residential (R) and non-residential (N) scenarios with varying instruction styles (Basic and Scene). Reorder: ✗ = LLM-guided Temporal Reordering disabled, ✔ = enabled. Router: Random = randomly select skill-based agents without utilizing action router; Qwen = Qwen2.5-VL-7B-Instruct; GLM = GLM-4.1V-9B-Thinking.
@@
Temporal Reordering Module. We conduct an ablation study to evaluate SkillNav’s two key components: the LLM-guided Temporal Reordering module and the VLM-based action router. The results, shown in Table 4, are reported across GSA-R2R splits, covering both residential (R) and non-residential (N) environments with varying instruction styles. First, we evaluate the effectiveness of the temporal reordering module. As shown in rows #2 and #4, when using the same router (Qwen2.5-VL-7B-Instruct), incorporating the reordering module consistently improves performance across all benchmarks. Notably, in Test-N-Basic, SPL increases $1.72\%$, demonstrating that temporally structured subgoals offer clearer guidance for effective skill selection.
@@
Action Router. To evaluate the effectiveness of our action router, we compare the performance of randomly selected skills without a router (row #1) against our proposed Qwen router. The observed improvements in both SR and SPL metrics clearly indicate the router’s effectiveness: specifically, Test-N-Scene SR increases from $54.61\%$ to $55.46\%$, and SPL rises notably from $43.17\%$ to $45.43\%$. These results confirm that our VLM-based router effectively selects appropriate skills even in the absence of temporal structuring. We further examine the significance of router selection by comparing rows #3 and #4, where the instruction reordering is fixed, and only the router model varies. Qwen2.5-VL-7B-Instruct consistently achieves superior SPL across all splits, particularly notable in Test-N-Scene ($47.96\%$ vs. $46.51\%$), underscoring its enhanced visual grounding capabilities compared to GLM-4.1V-9B-Thinking (Team et al. 2025). This emphasizes that high-quality vision-language representations are essential for effective skill routing.
Efficiency Analysis
Training Cost. Fine-tuning five skills on the Temporal Order Planning agent with R2R and synthetic skill-specific datasets requires approximately $3,329$ minutes ($\sim 55.5$ hours) in total. For comparison, SRDF training on R2R with larger data augmentation takes $2,521$ minutes ($\sim 42$ hours), suggesting that SkillNav’s skill-based training introduces a relatively higher training cost. However, this represents a one-time training investment; unlike prior supervised VLN models that require repeated retraining to adapt to new environments or instruction styles, SkillNav achieves strong generalization across datasets without additional retraining.
Table 5: Runtime and throughput of baselines and SkillNav. Numbers are wall-clock runtime in seconds. Random = randomly select skill-based agents without utilizing the action router.
Method Split Runtime (s) Inferences/s
Supervised VLN
ScaleVLN Test-R-Basic $513.8$ $\mathbf{28.03}$
Test-N-Basic $342.7$ $\mathbf{26.26}$
LLM-based VLN
MapGPT Test-R-Basic $\sim 597,000$ $0.02$
Test-N-Basic $\sim 373,000$ $0.02$
Our Mixture of Skill-based VLN
Random (ours) Test-R-Basic $2,223.4$ $6.48$
Test-N-Basic $1,507.9$ $5.97$
SkillNav (ours) Test-R-Basic $\sim 27,000$ $0.54$
Test-N-Basic $\sim 18,360$ $0.49$
Inference Cost. We provide inference time and throughput comparison in Table 5. SkillNav introduces overhead due to its Temporal Reordering LLM and VLM-based action router, reaching $0.49$ throughput on Test-N-Basic of GSA-R2R, which is roughly $50\times$ slower than ScaleVLN but still nearly $20\times$ faster than MapGPT. The Random variant, despite sharing the DUET as the backbone and selecting only one DUET for action prediction, is $4.3\times$ slower than ScaleVLN due to the per-observation skill selection overhead that prevents batch inference. Overall, while SkillNav is less efficient than supervised models, it achieves a better efficiency-generalization trade-off. Also, it advances both efficiency and generalization compared to LLM-based VLN agents.
@@
Figure 3 shows two qualitative examples highlighting SkillNav’s capability to dynamically select the appropriate skill at each navigation step. These examples illustrate the effectiveness of our approach in reordering temporal action plans, accurately identifying the currently focused subgoal via the router, and subsequently selecting the correct action. Specifically, in Figure 3 (a), the router correctly reasons that the agent has reached the target pillars and decides it is time to stop, resulting in the agent appropriately choosing the stop action at the view containing the pillars. Similarly, in Figure 3 (b), the router identifies the need to move toward the stairs and accordingly selects the vertical movement skill. Overall, SkillNav successfully interprets diverse instruction styles and performs robustly across both residential and non-residential scenes.
@@ Conclusion
We introduce SkillNav, a VLN agent that combines skill-based learning with VLM-based routing to dynamically select the most suitable actions based on the decision of the most relevant expert. We evaluate SkillNav on R2R to show strong navigation performance and demonstrate its generalization capabilities on the GSA-R2R dataset. While the utilization of LLM for temporal reordering and VLM for routing introduces computational overhead, SkillNav is more efficient than relying solely on LLMs or VLMs for navigation and achieves stronger performance than supervised VLN agents by exploiting both paradigms. Our framework provides a novel and interpretable approach that advances compositional reasoning and generalization for the VLN research community.
@@ Acknowledgment
Appendix A Primary Factors of Trajectory Generation
As introduced in Section Skill-Specific Data Synthesis and Agent Training in Methodology, we construct 5 skill-specific datasets and train the agents based on them. The primary factors for the construction of each skill are as follows:
@@ Appendix A Primary Factors of Trajectory Generation
(1) The stop should occur at a place with or after semantically relevant context for pausing, e.g., in front of a painting, at the foot of stairs. (2) The candidate image is very similar to the previous viewpoints.
@@
Appendix B Path Length in Trajectory Generation
@@
Appendix C Temporal Order Planning Agent
@@
Appendix D Stop and Pause Agent
Appendix E Efficiency Analysis
All experiments in efficiency analysis in Section Efficiency Analysis run on NVIDIA A6000. For the inference cost in Table 5, the number of predictions is $14,400$ for Test-R-Basic and $9,000$ for Test-N-Basic. For fairness, MapGPT is re-implemented with Qwen2.5-VL-7B-Instruct.
Appendix F LLM Usage
We used LLM-based tools for polishing grammar and aiding the writing. In addition, we utilize LLM to generate synthetic instructions for skill-specific datasets as described in Section Skill-Specific Data Synthesis and Agent Training. Moreover, LLMs and VLMs serve as our temporal reordering module and action router in Section Temporal Reordering Module and  VLM-based Action Router.
Appendix G Limitations
First, SkillNav is evaluated only in discrete VLN simulator environments (R2R, GSA-R2R, and NavNuances), leaving open the challenge of extending to continuous or real-world robotic navigation. Second, the approach depends on synthetic, skill-specific datasets generated via prompting, which may introduce distributional biases compared to human-authored instructions. We do a human evaluation on 20 cases with action routing, and the result shows 100% accuracy. This means with high confidence, the true accuracy is at least  84% on R2R Val Unseen.
Appendix H LLM and VLM Prompts
@@
2<Task>:
3Your task is to analyze a list of previously observed images and a natural language instruction.
4Determine which parts of the instruction have already been completed, and return the next step to be executed.
5<Response Rules>
6Your response must:
7- Return the next action using *exact phrasing* from the reordered instruction (no paraphrasing).
8- Match the sub-instruction to the visual context from previous images.
9- If the goal (e.g., pool table) has clearly been reached, return the final sub-instruction.
10- If *all* sub-instructions have been completed based on the visual path, do not return anything further. Stop reasoning.
11- If the final destination has been reached and the last step is a positional or waiting action (e.g., ‘‘wait there’’, ‘‘step to the left’’), return that as the next step.
12- You must reason about whether the agent is already at the destination.
13- If the current image shows the goal destination (e.g., inside the room with the pool table, or inside the open doorway), and the instruction contains a final step like ‘‘wait’’ or ‘‘adjust your position’’, that is the next sub-instruction.
14---
15Use the following reasoning strategy to determine what to do next:
16<Step-by-Step Reasoning Instructions>:
@@
18- Break the full instruction into smaller steps. Each sentence or clause typically represents one step.
19- Example:
@@
26- Do not reissue any previously executed sub-instructions.
27- Compare upcoming steps against what may have been visually completed, even if not explicitly executed one-by-one.
@@
29- Use visual context to infer if *multiple* sub-instructions have been completed in a single transition.
30- If image progression clearly shows the agent has already bypassed an intermediate area or reached a later goal, mark those steps as implicitly complete.
@@
32- If the current image shows the agent at or beyond the target of a sub-instruction, that step can be considered completed.
33- If the current image shows the agent inside the goal location and only a final positional instruction remains (e.g., ‘‘Step slightly to the left’’), return that final instruction.
@@
35- Use exact wording from the original instruction.
36- Do not return instructions that the agent already visually fulfilled, even if they were skipped.
@@
38{
39"Sub-instruction to be executed": "<exact next instruction clause>",
40"Reasoning": "<why this is the next step based on image sequence>"
41}
42CHECKPOINT:
43If multiple sub-instructions were completed based on a single or continuous image segment, skip them and jump to the next logical, visually unfulfilled step.
44---
@@
46Now, using the instruction and the visual history, identify the next step.
47IMPORTANT: Your response must be a valid JSON object without any surrounding text, code blocks, or explanations.
48Do not include markdown formatting like ‘‘‘json or ‘‘‘.
49
50<Original Whole Instruction>:
51‘‘{instruction}’’
52<Previous Sub-Instructions>:
53‘‘{previous_sub_instructions}’’
54<Previous Viewpoint Images>:
@@
The skill router determines which skill-based agent is most appropriate for executing the selected subgoal among the 5 skill-based agents. Besides, it receives the original instruction as contextual input to capture additional linguistic cues such as verbs and spatial references. It also uses the reasoning trace from the subgoal localizer to enhance its understanding of the current subgoal. The whole process is displayed in Listing  5.